OpenStudy (anonymous):

solve the equation. log(base 2)x+log(base 9)x=1

OpenStudy (anonymous):

\[\log_{2}X+\log_{9}X=1 \]

OpenStudy (anonymous):

There must be a nicer way but here's what I did. rearrange the logarithms into exponential form: \[\large 9^a =x\] \[\large 2^b = x\] which leads to the equation: \[\large a+b = 1\] but now we need one more equation since there are now three variables and only two equations. \[\large log_{2}x+\log_{9}x =1\] \[\huge 2^{\log_{2}x+\log_{9}x}= 2^1\] \[\huge 2^{\log_{2}x}*2^{\log_{9}x}=2\] \[\huge x*2^{\log_{9}x}=2\] \[\huge x=\frac{2}{2^{\log_9x}}\] sub from above: \[\large [\log_9x=a]\] \[\huge x=\frac{2}{2^{a}}\] \[\huge x = 2^{(1-a)}\]

OpenStudy (anonymous):

like wise if we raise everything to 9 in the initial step we arrive at: \[\huge x = 9(1-b)\] since x = x we get: \[\huge 2^{(1-a)}=9^{(1-b)}\] this with equation: \[\huge [a+b=1]\] allows you to solve either a or b, after which x is readily solvable

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

@ladyalice was able to clean it up a bit and make it more straight forward

OpenStudy (anonymous):

@Ladyalice25 You can ignore the other stuff, the most direct way consists of breaking the logs using the following log identity: \[\large \log_a b = \frac{\log_{10}b}{\log_{10}a}=\frac{\log b}{\log a}\] you can collect the x term within a few steps and the rest is straight algebra. \[\large \log_2x + \log_9x = 1\] \[\large \frac{\log x}{\log 2}+\frac{\log x}{\log 9}=1\] \[\large \log x(\frac{1}{\log2}+\frac{1}{\log9})=1\] \[\large = \log x(\frac{\log 9 +\log 2}{\log 2 * \log 9})=1\] \[\large \log x (\frac{\log 18}{\log 2 * \log 9})=1\] \[\large \log x = \frac{\log 2 * \log 9}{\log 18}\] \[\huge x = 10^{(\frac{\log 2 * \log 9}{\log 18})}\]